# DyGMamba: Efficiently Modeling Long-Term Temporal Dependency on Continuous-Time Dynamic Graphs with State Space Models

### v3

@@ Abstract
Learning useful representations for continuous-time dynamic graphs (CTDGs) is challenging, due to the concurrent need to span long node interaction histories and grasp nuanced temporal details. In particular, two problems emerge: (1) Encoding longer histories requires more computational resources, making it crucial for CTDG models to maintain low computational complexity to ensure efficiency; (2) Meanwhile, more powerful models are needed to identify and select the most critical temporal information within the extended context provided by longer histories. To address these problems, we propose a CTDG representation learning model named DyGMamba, originating from the popular Mamba state space model (SSM). DyGMamba first leverages a node-level SSM to encode the sequence of historical node interactions. Another time-level SSM is then employed to exploit the temporal patterns hidden in the historical graph, where its output is used to dynamically select the critical information from the interaction history. We validate DyGMamba experimentally on the dynamic link prediction task. The results show that our model achieves state-of-the-art in most cases. DyGMamba also maintains high efficiency in terms of computational resources, making it possible to capture long temporal dependencies with a limited computation budget.
@@ 1 Introduction
Dynamic graphs store node interactions in the form of links labeled with timestamps (Kazemi et al., 2020). In recent years, learning dynamic graphs has gained increasing interest since it can be used to facilitate various real-world applications. Dynamic graphs can be classified into two types, i.e., discrete-time dynamic graph (DTDG) and continuous-time dynamic graph (CTDG). A DTDG is represented as a sequence of graph snapshots that are observed at regular time intervals, where all the edges in a snapshot are taken as existing simultaneously, while a CTDG consists of a stream of events where each of them is observed individually with its own timestamp. Previous work (Kazemi et al., 2020) has indicated that CTDGs have an advantage over DTDGs in preserving temporal details, and therefore, more attention is paid to developing novel CTDG modeling approaches for dynamic graph representation learning.
Recent effort in CTDG modeling has resulted in a wide range of models. However, most of them are unable to model long-term temporal dependencies of nodes. To solve this problem, Yu et al. (2023) propose a CTDG model DyGFormer that can handle long-term node interaction histories based on Transformer (Vaswani et al., 2017). Despite its ability in modeling longer histories, employing a Transformer naturally introduces excessive usage of computational resources due to its quadratic complexity. Another recent work CTAN (Gravina et al., 2024) tries to capture long-term temporal dependencies by propagating graph information in a non-dissipative way over time with a graph convolution-based model. Despite the model’s high efficiency, Gravina et al. (2024) show that CTAN cannot capture very long histories and is surpassed by DyGFormer on the CTDGs where learning from very far away temporal information is critical. Based on these observations, we summarize the first challenge in CTDG modeling: How to develop a model that is scalable in modeling very long-term historical interactions? Another point worth noting is that as longer histories introduce more temporal information, more powerful models are needed to identify and select the most critical parts. This reveals another challenge: How to effectively select critical temporal information with long node interaction histories?
@@
To this end, we propose a new CTDG model named DyGMamba. DyGMamba first leverages a node-level Mamba SSM to encode historical node interactions. Another time-level Mamba SSM is then employed to exploit the edge-specific temporal patterns, where its output is used to dynamically select the critical information from the interaction history. To summarize: (1) We present DyGMamba, the first model using SSMs for CTDG representation learning; (2) DyGMamba demonstrates high efficiency and strong effectiveness in modeling long-term temporal dependencies in CTDGs; (3) Experimental results show that DyGMamba achieves new state-of-the-art on dynamic link prediction over most commonly-used CTDG datasets.
@@ 2.1 Related Work
Transformer (Vaswani et al., 2017) is a de facto backbone architecture in modern deep learning. However, its self-attention mechanism results in large space and time complexity, making it unsuitable for extremely long sequence modeling (Duman Keles et al., 2023). To address this, many works focus on building structured state space models that scale linearly or near-linearly with input sequence length (Gu et al., 2021; 2022b; 2022a; Smith et al., 2023; Peng et al., 2023; Ma et al., 2023; Gu & Dao, 2023). Most structured SSMs exhibit linear time invariance (LTI), meaning their parameters are not input-dependent and fixed for all time-steps. Gu & Dao (2023) demonstrate that LTI prevents SSMs from effectively selecting relevant information from the input context, which is problematic for tasks requiring context-aware reasoning. To solve this issue, Gu & Dao (2023) proposes S6, also known as Mamba, which uses a selection mechanism to dynamically choose important information from input sequence elements. Selection mechanism involves learning functions that map input data to SSM’s parameters, making Mamba both efficient and effective in modeling language, DNA sequences, and audio. Recently, there have been several works employing Mamba SSM for representation learning on static graphs (Wang et al., 2024; Behrouz & Hashemi, 2024) and spatial-temporal graphs (Li et al., 2024b). Unlike our work, they do not focus on capturing the dynamic evolution of graph structure and are thus not suitable for CTDG modeling.
@@ 2.2 Preliminaries
$\mathbf{p}\in\mathbb{R}^{|\mathbf{p}|}$ and $\mathbf{q}\in\mathbb{R}^{|\mathbf{p}|}$ are input and output sequences, where $|\mathbf{p}|$ is the sequence length of $\mathbf{p}$. * denotes the element-wise multiplication. When the dimension size of each element $p_{\tau}$ in $\mathbf{p}$ becomes higher (i.e., $p_{\tau}\in\mathbb{R}^{d_{2}}$ is a vector and $d_{2}>1$), both S4 and Mamba are in a Single-Input Single-Output (SISO) fashion, processing each input dimension in parallel with the same set of parameters. We follow Gu & Dao (2023) and denote the computation in Eq. 2 on the input sequences with vector elements as a function $\text{SSM}_{\bar{\mathbf{A}},\bar{\mathbf{B}},\mathbf{C}}(\cdot)$222 Input and output of $\text{SSM}_{\bar{\mathbf{A}},\bar{\mathbf{B}},\mathbf{C}}(\cdot)$ are matrices where each row is a vector corresponding to an element. See App. H for more details of SISO and the function.. Different from S4 which uses same parameters to process each element, Mamba changes its parameters into input-dependent by employing several trainable linear layers to map input into $\bar{\mathbf{B}}$, $\mathbf{C}$ and $\Delta$. The system is evolving as it processes different elements in the input sequence, making Mamba time-variant and suitable for modeling temporal sequences.
@@ 4 Experiments
In Sec. 4.2.1, we validate DyGMamba’s ability in CTDG representation learning by comparing it with baseline methods on dynamic link prediction [^3] To supplement, we also validate on the dynamic node classification task. Since current mainstream datasets of this task requires no long-term temporal reasoning, we put the discussion in App. F. Additionally, we also benchmark DyGMamba on DTDGs in App. L. This serves as supplementary experiment and does not directly connect to our main focus.. We show the effectiveness of model components by conducting ablation studies (Sec. 4.2.2) and analysis on synthetic datasets (Sec. 4.2.3). In Sec. 4.3.1, we show DyGMamba’s efficiency against various baselines. We also show that it achieves much stronger scalability in modeling long-term temporal information compared with the current state-of-the-art DyGFormer (Sec. 4.3.2 and Sec. 4.3.3).
@@ 4.2.2 Ablation Study
We conduct four ablation studies to study the effectiveness of model components. In study A, we make a model variant (Variant A) by removing the time-level SSM block and restrain our model from learning temporal patterns (information selection is substituted by mean pooling over the output of Eq. 4). In study B, we make a model variant (Variant B) by removing the Mamba SSM layers (Eq. 3) in the node-level SSM block. In study C, we switch the computation of the selection weights $\beta_{\theta}$ in Eq. 6d to $\beta_{\theta}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{\theta}))$ ($f_{\text{sel}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$) to create Variant C. In study D, we base on Variant C and develop Variant D that further enables information selection from opposite nodes, i.e., $\beta_{u}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{v}))$/ $\beta_{v}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{u}))$. Both ablation C and D do information selection without learning temporal patterns. From Table 3 and 4, we find that: (1) Variant A is constantly beaten by DyGMamba, showing the effectiveness of dynamic information selection based on edge-specific temporal patterns; (2) DyGMamba always outperforms Variant B, indicating the importance of encoding the one-hop temporal neighbors with SSM layers for capturing graph dynamics; (3) Variant C and D perform better than Variant A in most cases, implying that selecting temporal information is generally contributive; (4) Variant C generally lags behind Variant D, meaning that information selection from opposite node is beneficial; (5) DyGMamba performs better than both Variant C and D in almost all cases, proving that information selection based on temporal patterns is more effective.
@@ 4.3.1 Model Size, Per Epoch Training Time and GPU Memory Comparison Across Various Models
Fig. 2 compares DyGMamba with five baselines in terms of number of parameters (model size), per epoch training time, and GPU memory consumption during training [^4] The baselines not included here are either extremely inefficient (e.g., CAWN) or inferior in performance (e.g., DyRep). Complete statistics of all baseline models presented in App. G. We find that: (1) DyGMamba uses very few parameters while maintaining the best performance, showing a strong parameter efficiency. Only CTAN constantly uses fewer parameters than DyGMamba, however, its performance is significantly worse with the only exception of Enron; (2) DyGMamba is always more efficient than DyGFormer with the same length of input sequence ($\rho/p=$32) while achieving the same performance; (3) Although DyGMamba generally consumes more GPU memory and takes more time to train per epoch compared with most baselines, the gap of consumption is modest. To model more temporal neighbors for long-range temporal dependent datasets, DyGMamba naturally requires more computational resources, thus enlarging the consumption gap. DyGFormer shows the same trend as DyGMamba since it also captures long-term temporal dependencies but at a higher cost; (4) CTAN requires very few computational resources. However, on long-range temporal dependent datasets, it is beaten by DyGFormer and DyGMamba by a large margin, e.g., on LastFM and Enron. Besides, CTAN is also hard to converge. Although it takes little time to train a single epoch, it needs more epochs to reach the best performance, leading to a long total training time. See App. G.2 for a total training time comparison among DyGFormer, CTAN and DyGMamba. To supplement, we also present in App. G.3 another experiment to compare DyGMamba with baselines in modeling an increasing number of temporal neighbors on Enron with limited total training time.
Figure 2: Efficiency comparison on four datasets among DyGMamba and five baselines in terms of number (#) of parameters, training time per epoch and GPU memory. The performance metric here is AP of transductive link prediction under random NSS. The greener, the better performance/efficiency. In contrast to other methods, DyGMamba consistently shows strong overall capability across different datasets. More explanations in Sec. 4.3.1.

### v4

@@ Abstract
Learning useful representations for continuous-time dynamic graphs (CTDGs) is challenging, due to the concurrent need to span long node interaction histories and grasp nuanced temporal details. In particular, two problems emerge: (1) Encoding longer histories requires more computational resources, making it crucial for CTDG models to maintain low computational complexity to ensure efficiency; (2) Meanwhile, more powerful models are needed to identify and select the most critical temporal information within the extended context provided by longer histories. To address these problems, we propose a CTDG representation learning model named DyGMamba, originating from the popular Mamba state space model (SSM). DyGMamba first leverages a node-level SSM to encode the sequence of historical node interactions. Another time-level SSM is then employed to exploit the temporal patterns hidden in the historical graph, where its output is used to dynamically select the critical information from the interaction history. We validate DyGMamba experimentally on the dynamic link prediction task. The results show that our model achieves state-of-the-art in most cases. DyGMamba also maintains high efficiency in terms of computational resources, making it possible to capture long temporal dependencies with a limited computation budget [^1] We release our code at the following link: https://github.com/ZifengDing/DyGMamba..
@@ 1 Introduction
Dynamic graphs store node interactions in the form of links labeled with timestamps (Kazemi et al., 2020). In recent years, learning dynamic graphs has gained increasing interest since it can be used to facilitate various real-world applications. Dynamic graphs can be classified into two types, i.e., discrete-time dynamic graph (DTDG) and continuous-time dynamic graph (CTDG). A DTDG is represented as a sequence of graph snapshots that are observed at regular time intervals, where all the edges in a snapshot are taken as existing simultaneously, while a CTDG consists of a stream of events where each of them is observed individually with its own timestamp. Previous works (Kazemi et al., 2020; Shirzadkhani et al., 2024) have indicated that CTDGs have an advantage over DTDGs in preserving temporal details, and therefore, more attention is paid to developing novel CTDG modeling approaches for dynamic graph representation learning.
Recent effort in CTDG modeling has resulted in a wide range of models. However, most of them are unable to model long-term temporal dependencies of nodes, despite the existence of abundant historical information. To solve this problem, Yu et al. (2023) propose a CTDG model DyGFormer that can handle long-term node interaction histories based on Transformer (Vaswani et al., 2017). Despite its ability in modeling longer histories, employing a Transformer naturally introduces excessive usage of computational resources due to its quadratic complexity. Another recent work CTAN (Gravina et al., 2024) tries to capture long-term temporal dependencies by propagating graph information in a non-dissipative way over time with a graph convolution-based model. Despite the model’s high efficiency, Gravina et al. (2024) show that CTAN cannot capture very long histories and is surpassed by DyGFormer on the CTDGs where learning from very far away temporal information is critical. Based on these observations, we summarize the first challenge in CTDG modeling: How to develop a model that is scalable in modeling very long-term historical interactions? Another point worth noting is that as longer histories introduce more temporal information, more powerful models are needed to identify and select the most critical parts. This reveals another challenge: How to effectively select critical temporal information with long node interaction histories?
@@
To this end, we propose a new CTDG model named DyGMamba. DyGMamba first leverages a node-level Mamba SSM to encode historical node interactions. Another time-level Mamba SSM is then employed to exploit the edge-specific temporal patterns, where its output is used to dynamically select the critical information from the interaction history. To summarize: (1) We present DyGMamba, the first model using SSMs for CTDG representation learning; (2) DyGMamba demonstrates high efficiency and strong effectiveness in modeling long-term temporal dependencies in CTDGs; (3) Experimental results show that DyGMamba achieves new state-of-the-art on dynamic link prediction over most common CTDG datasets.
@@ 2.1 Related Work
Transformer (Vaswani et al., 2017) is a de facto backbone architecture in modern deep learning. However, its self-attention mechanism results in large space and time complexity, making it unsuitable for extremely long sequence modeling (Duman Keles et al., 2023). To address this, many works focus on building structured state space models that scale linearly or near-linearly with input sequence length (Gu et al., 2021; 2022b; 2022a; Smith et al., 2023; Peng et al., 2023; Ma et al., 2023; Gu & Dao, 2023). Most structured SSMs exhibit linear time invariance (LTI), meaning their parameters are not input-dependent and fixed for all time-steps. Gu & Dao (2023) demonstrate that LTI prevents SSMs from effectively selecting relevant information from the input context, which is problematic for tasks requiring context-aware reasoning. To solve this issue, Gu & Dao (2023) proposes S6, also known as Mamba, which uses a selection mechanism to dynamically choose important information from input sequence elements. Selection mechanism involves learning functions that map input data to SSM’s parameters, making Mamba both efficient and effective in modeling language, DNA sequences, and audio.
State Space Models for Graphs.
Recently, there have been several works employing Mamba SSM for representation learning on static graphs (Wang et al., 2024; Behrouz & Hashemi, 2024). Wang et al. (2024) integrate a Mamba SSM into graph transformers, employing input-dependent node selection and prioritization mechanisms to efficiently capture long-range multi-hop dependencies in static graphs. Behrouz & Hashemi (2024) adapt Mamba by creating a bidirectional SSM for static graphs, utilizing hierarchical tokenization and an optional positional encoding mechanism to improve scalability and performance on diverse graph-based downstream tasks. Both of them aim to efficiently encode long-range multi-hop graph information. Unlike our work, they do not have dedicated structures to process temporal information and are thus not suitable for CTDG modeling. There is also another work (Li et al., 2024b) that tries to use Mamba to model spatial-temporal graphs (STGs). Li et al. (2024b) employ a Spatial-Temporal Selective State Space Module for adaptive spatial-temporal feature selection and Kalman Filtering Graph Neural Networks to dynamically integrate and optimize embeddings from various temporal granularities. Although this work achieves high efficiency and superior forecasting accuracy on STGs, it is not applicable to modeling CTDGs because (1) STGs are inherently DTDGs, and (2) this model primarily aims at predicting node states in STGs (e.g., traffic sensor readings) rather than explicitly capturing the continuous and irregular structural evolution characteristic of CTDGs. In our work, we propose a Mamba-based model to capture evolving graph dynamics by encoding long-range historical event streams and introducing a specialized module designed explicitly to encode edge-specific temporal patterns. This approach significantly differs from existing Mamba-based graph reasoning methods, enabling more effective temporal reasoning on CTDGs.
@@ 2.2 Preliminaries
$\mathbf{p}\in\mathbb{R}^{|\mathbf{p}|}$ and $\mathbf{q}\in\mathbb{R}^{|\mathbf{p}|}$ are input and output sequences, where $|\mathbf{p}|$ is the sequence length of $\mathbf{p}$. * denotes the element-wise multiplication. When the dimension size of each element $p_{\tau}$ in $\mathbf{p}$ becomes higher (i.e., $p_{\tau}\in\mathbb{R}^{d_{2}}$ is a vector and $d_{2}>1$), both S4 and Mamba are in a Single-Input Single-Output (SISO) fashion, processing each input dimension in parallel with the same set of parameters. We follow Gu & Dao (2023) and denote the computation in Eq. 2 on the input sequences with vector elements as a function $\text{SSM}_{\bar{\mathbf{A}},\bar{\mathbf{B}},\mathbf{C}}(\cdot)$333 Input and output of $\text{SSM}_{\bar{\mathbf{A}},\bar{\mathbf{B}},\mathbf{C}}(\cdot)$ are matrices where each row is a vector corresponding to an element. See App. H for more details of SISO and the function.. Different from S4 which uses same parameters to process each element, Mamba changes its parameters into input-dependent by employing several trainable linear layers to map input into $\bar{\mathbf{B}}$, $\mathbf{C}$ and $\Delta$. The system is evolving as it processes different elements in the input sequence, making Mamba time-variant and suitable for modeling temporal sequences.
@@ 4 Experiments
In Sec. 4.2.1, we validate DyGMamba’s ability in CTDG representation learning by comparing it with baseline methods on dynamic link prediction [^4] To supplement, we also validate on the dynamic node classification task. Since current mainstream datasets of this task requires no long-term temporal reasoning, we put the discussion in App. F. Additionally, we also benchmark DyGMamba on DTDGs in App. K. This serves as supplementary experiment and does not directly connect to our main focus.. We show the effectiveness of model components by conducting ablation studies (Sec. 4.2.2) and analysis on synthetic datasets (Sec. 4.2.3). In Sec. 4.3.1, we show DyGMamba’s efficiency against various baselines. We also show that it achieves much stronger scalability in modeling long-term temporal information compared with the current state-of-the-art DyGFormer (Sec. 4.3.2 and Sec. 4.3.3).
@@ 4.2.2 Ablation Study
We conduct four ablation studies to study the effectiveness of model components. In study A, we make a model variant (Variant A) by removing the time-level SSM block and restrain our model from learning temporal patterns (information selection is substituted by mean pooling over the output of Eq. 4). In study B, we make a model variant (Variant B) by removing the Mamba SSM layers (Eq. 3) in the node-level SSM block. In study C, we switch the computation of the selection weights $\beta_{\theta}$ in Eq. 6d to $\beta_{\theta}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{\theta}))$ ($f_{\text{sel}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$) to create Variant C. In study D, we base on Variant C and develop Variant D that further enables information selection from opposite nodes, i.e., $\beta_{u}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{v}))$/ $\beta_{v}=\text{Softmax}(f_{\text{sel}}(\mathbf{H}^{t}_{u}))$. Both ablation C and D do information selection without learning temporal patterns. In study E and F, we devise Variant E and Variant F by removing the time features and the node interaction frequency features, respectively, during neighbor encoding. From Table 3 and 4, we find that: (1) Variant A is constantly beaten by DyGMamba, showing the effectiveness of dynamic information selection based on edge-specific temporal patterns; (2) DyGMamba always outperforms Variant B, indicating the importance of encoding the one-hop temporal neighbors with SSM layers for capturing graph dynamics; (3) Variant C and D perform better than Variant A in most cases, implying that selecting temporal information is generally contributive; (4) Variant C generally lags behind Variant D, meaning that information selection from opposite node is beneficial; (5) DyGMamba performs better than both Variant C and D in almost all cases, proving that information selection based on temporal patterns is more effective; (6) Variant E and F in general achieve worse results than DyGMamba, validating the contribution of both time and node interaction frequency features.
@@ 4.3.1 Model Size, Per Epoch Training Time and GPU Memory Comparison Across Various Models
Fig. 2 compares DyGMamba with five baselines in terms of number of parameters (model size), per epoch training time, and GPU memory consumption during training [^5] The baselines not included here are either extremely inefficient (e.g., CAWN) or inferior in performance (e.g., DyRep). Complete statistics of all baseline models presented in App. G.1. We find that: (1) DyGMamba uses very few parameters while maintaining the best performance, showing a strong parameter efficiency. Only CTAN constantly uses fewer parameters than DyGMamba, however, its performance is significantly worse with the only exception of Enron; (2) DyGMamba is always more efficient than DyGFormer with the same length of input sequence ($\rho/p=$32) while achieving the same performance; (3) Although DyGMamba generally consumes more GPU memory and takes more time to train per epoch compared with most baselines, the gap of consumption is modest. To model more temporal neighbors for long-range temporal dependent datasets, DyGMamba naturally requires more computational resources, thus enlarging the consumption gap. DyGFormer shows the same trend as DyGMamba since it also captures long-term temporal dependencies but at a higher cost; (4) CTAN requires very few computational resources. However, on long-range temporal dependent datasets, it is beaten by DyGFormer and DyGMamba by a large margin, e.g., on LastFM and Enron. Besides, CTAN is also hard to converge. Although it takes little time to train a single epoch, it needs more epochs to reach the best performance, leading to a long total training time. See App. G.2 for a total training time comparison among DyGFormer, CTAN and DyGMamba. To supplement, we also present in App. G.3 another experiment to compare DyGMamba with baselines in modeling an increasing number of temporal neighbors on Enron with limited total training time.
Figure 2: Efficiency comparison on four datasets among DyGMamba and five baselines in terms of number (#) of parameters, training time per epoch and GPU memory. The performance metric here is AP of transductive link prediction under random NSS. The greener, the better performance/efficiency. In contrast to other methods, DyGMamba consistently shows strong overall capability across different datasets. More explanations in Sec. 4.3.1. Further comparison on Reddit and Social Evo. is presented in App. G.1
@@ 4.3.2 Impact of Patch Size on Scalability and Performance
5 Discussion
In this section, we discuss where DyGMamba excels and where it falls short, and explain our choice to consider only one-hop temporal neighbors. We additionally provide a full justification of our motivation in App. L, recommended for reading after the main paper.
Where DyGMamba Excels and Where It Falls Short.
We summarize here the conditions under which DyGMamba performs well and scenarios where it may fall short. DyGMamba is expected to outperform previous state-of-the-art models in the following scenarios: (1) when modeling long-range temporal information is critical; (2) when there is a clear temporal pattern, i.e., the density of repeated edges follows a specific trend. For the first scenario, we have demonstrated in Table 1 and 2 that DyGMamba surpasses previous models on long-range temporal-dependent datasets (LastFM, MOOC, and Enron) in both transductive and inductive link prediction under the random NSS setting. Under the historical and inductive NSS settings, DyGMamba consistently ranks among the top 3 models and even achieves the top 1 position in 7 out of 9 cases. Regarding the second scenario, Table 5 confirms that DyGMamba’s dynamic information selection module significantly enhances its capability to model temporal patterns, enabling superior performance on datasets characterized by clear temporal trends. Conversely, DyGMamba might underperform previous state-of-the-art models on certain datasets (e.g., Wikipedia and UCI) under historical and inductive NSS settings, particularly when the number of available one-hop temporal neighbors for nodes in the link prediction query is limited (the availability of historical neighbors across datasets for both link prediction tasks under all NSS settings is presented in Table 22). DyGMamba rely on abundant historical information to effectively capture graph dynamics and temporal patterns. Insufficient historical context may hinder its performance, leading to potential misclassification of negative edges and higher false positive rates.
Why We Only Consider One-Hop Neighbors.
We only consider encoding one-hop temporal neighbors in DyGMamba due to two reasons. First, our primary motivation is to design an efficient model that effectively captures long-range temporal dependencies in CTDGs. While incorporating higher-order information is an intuitive approach, it introduces several challenges. Maintaining the same computational budget would require reducing the number of first-order (one-hop) temporal neighbors, leading to a less representative one-hop neighborhood. The alternative—allocating more computational resources such as increased GPU memory—goes against our objective of efficiency in this work. Moreover, modeling higher-order temporal neighbors requires additional specialized modules to differentiate the influence of neighbors from different hops, which can further increase computational cost and difficulties in modeling. Besides, recent CTDG models that only consider one-hop temporal neighbors, such as DyGFormer and FreeDyG, have demonstrated strong performance, and DyGMamba further validates this observation. Some of our baselines, such as CAWN and CTAN, do incorporate higher-order information. However, as shown in Table 1 and 2, they are outperformed by DyGMamba in most cases. A potential reason for this is that these models may still struggle to optimally differentiate information from different hops. Efficiently integrating higher-order information into CTDG modeling is a promising research direction, and we would be happy to explore it in future work.